which equation gives the directrix of y-1/12x^2-5?
y=-8
y=3
x=3
y=-5
please need help for math class.

Answer by lwsshak3(11628)

which equation gives the directrix of y-1/12x^2-5?
y=-8
y=3
x=3
y=-5
**
I think you meant to write the equation as:y=(1/12)x^2-5
Which can be rewritten as: (1/12)x^2=y+5 or x^2=12(y+5)
This equation is of the standard form: (x-h)^2=4py, with (h,k) being the (x,y) coordinates of the vertex.
..
For given equation:
vertex: (0,-5)
Axis of symmetry: x=0 or y-axis (parabola opens upwards)
4p=12
p=3
The directrix is p units below the vertex on the axis of symmetry
Equation of Directrix: y= -8